OpenStudy (anonymous):

Which of the following is equal to sqrt(18x^9y^4)

OpenStudy (anonymous):

I think its 3x^3y^2sqrt( 2)

OpenStudy (anonymous):

like this? \[\sqrt{18x^{9y^4}}\]

OpenStudy (anonymous):

Like this \(\sqrt{18x^9 y^4}\)

OpenStudy (anonymous):

ok! ^_^

OpenStudy (anonymous):

so, x^9 = (x^2)(x^2)(x^2)(x^2)(x) and if I do sqrt ( (x^2)(x^2)(x^2)(x^2)(x) ) I get?

OpenStudy (anonymous):

I gotta leave soon, are you still here?

OpenStudy (anonymous):

there?

OpenStudy (anonymous):

\[sqrt(x^9)\]

OpenStudy (anonymous):

sorry about that i have to keep refreshing to see what you type :p

OpenStudy (anonymous):

it can equal that, it can also equal: sqrt ( (x^2)(x^2)(x^2)(x^2)(x) ) = x^4sqrt(x)

OpenStudy (anonymous):

see, sqrt(x^2) = x if we have four x^2 under the square root, that is the same as having 4 x's outside the square root \[ \sqrt{(x^2)(x^2)(x^2)(x^2)(x)} = x(x)x(x)\sqrt{x} = x^4\sqrt{x}\]

OpenStudy (anonymous):

so lets have you try the same thing with this: \[\sqrt{y^4} = \sqrt{y^?y^?} = y^?\]

OpenStudy (anonymous):

I see I get it ! so y^2 y^2 = y^4

OpenStudy (anonymous):

very good! make sure u dont' forget the square roots! \[\sqrt{y^4} = \sqrt{y^2y^2} = y^2\]

OpenStudy (anonymous):

so we have taken care of the y's and the x's, all that is left to look at is the 18 so this is where we are to then: \[y(x^4)\sqrt{18x} \]

OpenStudy (anonymous):

\[\sqrt{18x} = \sqrt{18}\sqrt{x}=\sqrt{2*9}\sqrt{x}=\sqrt{2}\sqrt{9}\sqrt{x}\]

OpenStudy (anonymous):

\[3 \sqrt(2) \sqrt(x^9 y^4)\] Right?

OpenStudy (anonymous):

that is a correct statment, but remember, we can reduce the sqrt(x^9y^4) part like we did before

OpenStudy (anonymous):

I gotta go, i'm sorry! you've done great!

OpenStudy (anonymous):

Ok thanks!